# The Disparate Benefits of Deep Ensembles

### v1

@@ 5 The Disparate Benefits Effect of Deep Ensembles
Table 1: Disparate Benefits: Change in performance and fairness violations due to ensembling. Significant differences ($\Delta$) between the Deep Ensemble (c.f. Tab 2) and the average ensemble member (c.f. Tab. 3) are highlighted in bold (t-test, five runs, $p<0.05$). Gray cells denote that fairness violations are $>0.05$ for both the Deep Ensemble and the average ensemble member.
@@
FF age / gender $\bm{0.022_{\pm 0.001}}$-$\bm{0.022_{\pm 0.003}}$-$\bm{0.017_{\pm 0.004}}$-$\bm{0.017_{\pm 0.003}}$
FF age / race $\bm{0.022_{\pm 0.001}}$-$\bm{0.009_{\pm 0.003}}$-$\bm{0.012_{\pm 0.004}}$-$\bm{0.007_{\pm 0.003}}$
FF gender / age $\bm{0.014_{\pm 0.001}}$ $-0.001_{\pm 0.001}$ $\bm{-0.007_{\pm 0.001}}$ $\bm{-0.004_{\pm 0.002}}$
FF gender / race $\bm{0.014_{\pm 0.001}}$ $-0.001_{\pm 0.001}$-$0.000_{\pm 0.000}$ $-0.002_{\pm 0.002}$
FF race / age $\bm{0.015_{\pm 0.001}}$ $\bm{-0.004_{\pm 0.001}}$-$\bm{0.005_{\pm 0.002}}$ $\bm{-0.001_{\pm 0.000}}$
FF race / gender $\bm{0.015_{\pm 0.001}}$-$0.000_{\pm 0.002}$ $-0.008_{\pm 0.006}$-$0.002_{\pm 0.004}$
UTK age / gender $\bm{0.015_{\pm 0.001}}$-$\bm{0.017_{\pm 0.001}}$-$\bm{0.015_{\pm 0.002}}$-$\bm{0.012_{\pm 0.001}}$
UTK age / race $\bm{0.015_{\pm 0.001}}$-$\bm{0.010_{\pm 0.002}}$-$\bm{0.010_{\pm 0.001}}$-$\bm{0.004_{\pm 0.002}}$
UTK gender / age $\bm{0.009_{\pm 0.001}}$-$0.001_{\pm 0.001}$ $\bm{-0.006_{\pm 0.002}}$ $\bm{-0.003_{\pm 0.001}}$
UTK gender / race $\bm{0.009_{\pm 0.001}}$-$0.000_{\pm 0.001}$-$0.001_{\pm 0.002}$-$0.001_{\pm 0.001}$
UTK race / age $\bm{0.021_{\pm 0.001}}$-$\bm{0.013_{\pm 0.001}}$-$\bm{0.007_{\pm 0.002}}$-$0.000_{\pm 0.001}$
UTK race / gender $\bm{0.021_{\pm 0.001}}$-$0.003_{\pm 0.002}$ $-0.002_{\pm 0.003}$ $-0.002_{\pm 0.002}$
@@
CX age $\bm{0.005_{\pm 0.000}}$-$\bm{0.001_{\pm 0.000}}$-$\bm{0.008_{\pm 0.004}}$-$\bm{0.003_{\pm 0.001}}$
CX gender $\bm{0.005_{\pm 0.000}}$-$0.000_{\pm 0.001}$-$0.001_{\pm 0.004}$-$0.001_{\pm 0.002}$
CX race $\bm{0.005_{\pm 0.000}}$ $\bm{-0.002_{\pm 0.001}}$-$0.000_{\pm 0.003}$ $-0.001_{\pm 0.002}$
@@ 6 What is the Reason for Disparate Benefits?
In this section, we investigate the potential causes behind the disparate benefits effect. We first investigate how the per-group PR, TPR and FPR metrics change when adding ensemble members, as the considered fairness metrics (Eq. (5) - Eq. (7)) are derived from them. Although this provides insight about why the disparate benefits effect occurs, it lacks an explanation for the underlying cause. We hypothesize that the disparate benefits effect results from the predictive diversity among ensemble members. Our empirical results agree with this hypothesis, suggesting that a gap in average predictive diversity between groups is causing the disparate benefits effect. We conclude with a synthetic experiment to demonstrate the soundness of our hypothesis in a controlled setting.
@@
Figure 3: Average predictive diversity ($\overline{\text{DIV}}$) per group $A$ and target $Y$. Exemplary results for datasets FF, UTK and CX. Arrows indicate per-target group differences. Top row (a)–(c): Significant disparate benefits (c.f. Tab. 1) occur when $\overline{\text{DIV}}$ differences between groups are large. Bottom row (d)–(f): No significant disparate benefits occur when $\overline{\text{DIV}}$ differences are small.
@@
Changes to predictions for increasing ensemble size. We begin by examining how the metrics PR, TPR, and FPR for each group change when ensemble members are added, since the considered fairness metrics (Eq. (5) - Eq. (7)) are based on these. Fig. 3 shows these changes for the model trained on FF with age as target variable and gender as protected group, evaluated on the FF test dataset. The results show that the increase in SPD comes from a decrease in the PR of the disadvantaged group when adding ensemble members, while the PR of the advantaged group remains stable. The TPR of the disadvantaged group stays constant, but the TPR of the advantaged group increases, so the Deep Ensemble improves in correctly predicting $Y=1$ only for the advantaged group, resulting in a higher EOD. The FPR of both groups decreases, more so for the disadvantaged group, thus the Deep Ensemble improves in correctly predicting $Y=0$ (as FPR is one minus the true negative rate). However, this doesn’t offset the TPR disparity, resulting in higher AOD. Results for the remaining tasks are provided in Fig. 11 - Fig. 13 in the appendix.
@@
Predictive diversity of ensemble members. The ensemble predictive distribution (Eq. (1)) is an average over the predictive distributions of its members. Therefore, the origin of the disparate benefits effect must be in the characteristics of the predictive distributions of individual members. Previous work investigated the predictive diversity of individual members as the driving mechanism for the increase in the performance of Deep Ensembles (Abe et al., 2022b; Jeffares et al., 2023; Abe et al., 2024). Only if individual members have different predictive distributions, combining them can lead to an ensemble that performs better than the individual models. While previous work investigates predictive diversity for individual inputs $\bm{x}$, we are interested in the average predictive diversity on the test dataset $\mathcal{D}^{\prime}=\{(\bm{x}_{k},y_{k},a_{k})\}_{k=1}^{K}$. Following from the definition of predictive diversity by Jeffares et al. (2023) (Theorem 4.3), the average predictive diversity $\overline{\text{DIV}}$ is thus given by
@@
$\overline{\text{DIV}}=\frac{1}{K}\sum_{k=1}^{K}\underbrace{\log\left(\frac{1}{N}\sum_{n=1}^{N}p(y=y_{k}\mid\bm{x}_{k},\bm{w}_{n})\right)}_{\text{Ensemble Log-Likelihood}}\ -\ \underbrace{\frac{1}{N}\sum_{n=1}^{N}\log p(y=y_{k}\mid\bm{x}_{k},\bm{w}_{n})}_{\text{Average Member Log-Likelihood}}\ ,$ (8)
@@
for a test dataset $\mathcal{D}^{\prime}$, and a set of $N$ models with parameters $\{\bm{w}_{n}\}_{n=1}^{N}$. In Sec. C in the appendix we provide a more detailed discussion about the average predictive diversity and how it arises as a natural measure of interest from a Bayesian perspective on Deep Ensembles. In summary, the average predictive diversity estimates the ratio of the marginal likelihood of the test dataset (how well individual models agree with $\mathcal{D}^{\prime}$) and the likelihood under the Bayesian model average (how well the ensemble agrees with $\mathcal{D}^{\prime}$).
@@
We hypothesize that differences in the average predictive diversity per group cause the disparate benefits effect. To investigate this hypothesis, we consider two sets of tasks for FF, UTK and CX, respectively: those where the disparate benefits effect occurs and those where it does not occur (c.f. Tab. 1). The results are depicted in Fig. 3, showing the average predictive diversity $\overline{\text{DIV}}$ per combination of the target variable $Y$ and the protected group attribute $A$. In agreement with our hypothesis, tasks showing the disparate benefits effect (Fig. 3a-c) have substantial differences in average predictive diversity between groups, while tasks without the effect (Fig. 3d-f) show only minimal differences. Results on all tasks are given in Fig. 14 - Fig. 16 in the appendix.
(a) Change in performance and fairness for increasing sizes of the Deep Ensemble.
(b) Change in PR, TPR and FPR for increasing sizes of the Deep Ensemble.
@@
Figure 4: Controlled experiment. Top row: The performance (accuracy) increases whereas fairness (1-SPD, 1-EOD, 1-AOD) decreases when adding more members to the ensemble. Bottom row: The disparate benefits effect is caused by increased PR and TPR, as well as decreased FPR for the group with higher average predictive diversity $A=1$. For the group with smaller average predictive diversity $A=0$, there are no significant changes in PR, TPR and FPR.
Figure 5: Inputs per target and group.
@@
Controlled experiment. To test our hypothesis of the per-group differences in average predictive diversity $\overline{\text{DIV}}$ causing the disparate benefits effect, we conduct a controlled experiment. We use the FashionMNIST (Xiao et al., 2017) dataset and create a binary classification problem with two targets: “T-shirt/top” ($Y=0$) vs “Shirt” ($Y=1$), and two groups, $A=0$ where the same image of the same target is concatenated twice and $A=1$ where two different images of the same target are concatenated. This is done for both the train and test datasets. An illustration of inputs $\bm{x}$ for both targets and groups is given in Fig. 5. Naturally, having an input consisting of two different images ($A=1$) should lead to more diverse ensemble members, as they may learn to use the top image, the bottom image or any combination of features from both. The combination of two identical images ($A=0$) does not provide additional information and therefore should not lead to an increased diversity of the ensemble members. This intuition is experimentally confirmed by having a higher $\overline{\text{DIV}}$ for $A=1$ (Fig. 4c). We observe the same behavior regarding the change in performance, fairness (Fig. 4a) as well as PR, TPR and FPR (Fig. 4b) as for the real-world datasets we investigate throughout the rest of the paper. In sum, the synthetic dataset (Fig. 5) enforces more predictive diversity for one group (Fig. 4c), leading to the disparate benefits effect (Fig. 4a, b).
7 Mitigating the Negative Impact of Disparate Benefits

### v2

@@ 5 The Disparate Benefits of Deep Ensembles
Table 1: Disparate Benefits: Change in performance and fairness violations due to ensembling. Significant differences ($\Delta$) between the Deep Ensemble (cf. Tab 3) and the average ensemble member (cf. Tab. 4) are highlighted in bold (t-test, five runs, $p<0.05$). Gray cells denote that fairness violations are $>0.05$ for both the Deep Ensemble and the average ensemble member.
@@
FF age / gender .$\bm{.022_{\pm.001}}$.$\bm{.022_{\pm.003}}$.$\bm{.017_{\pm.004}}$.$\bm{.017_{\pm.003}}$
FF age / race .$\bm{.022_{\pm.001}}$.$\bm{.009_{\pm.003}}$.$\bm{.012_{\pm.004}}$.$\bm{.007_{\pm.003}}$
FF gender / age .$\bm{.014_{\pm.001}}$.$\text{-}.001_{\pm.001}- .$.$\bm{\text{-}.007_{\pm.001}}$.$\bm{\text{-}.004_{\pm.002}}$
FF gender / race .$\bm{.014_{\pm.001}}$.$\text{-}.001_{\pm.001}- .$.$000_{\pm.000}.$.$\text{-}.002_{\pm.002}- .$
FF race / age .$\bm{.015_{\pm.001}}$.$\bm{\text{-}.004_{\pm.001}}$.$\bm{.005_{\pm.002}}$.$\bm{\text{-}.001_{\pm.000}}$
FF race / gender .$\bm{.015_{\pm.001}}$.$000_{\pm.002}.$.$\text{-}.008_{\pm.006}- .$.$002_{\pm.004}.$
UTK age / gender .$\bm{.015_{\pm.001}}$.$\bm{.017_{\pm.001}}$.$\bm{.015_{\pm.002}}$.$\bm{.012_{\pm.001}}$
UTK age / race .$\bm{.015_{\pm.001}}$.$\bm{.010_{\pm.002}}$.$\bm{.010_{\pm.001}}$.$\bm{.004_{\pm.002}}$
UTK gender / age .$\bm{.009_{\pm.001}}$.$001_{\pm.001}.$.$\bm{\text{-}.006_{\pm.002}}$.$\bm{\text{-}.003_{\pm.001}}$
UTK gender / race .$\bm{.009_{\pm.001}}$.$000_{\pm.001}.$.$001_{\pm.002}.$.$001_{\pm.001}.$
UTK race / age .$\bm{.021_{\pm.001}}$.$\bm{.013_{\pm.001}}$.$\bm{.007_{\pm.002}}$.$000_{\pm.001}.$
UTK race / gender .$\bm{.021_{\pm.001}}$.$003_{\pm.002}.$.$\text{-}.002_{\pm.003}- .$.$\text{-}.002_{\pm.002}- .$
@@
CX age .$\bm{.005_{\pm.000}}$.$\bm{.001_{\pm.000}}$.$\bm{.008_{\pm.004}}$.$\bm{.003_{\pm.001}}$
CX gender .$\bm{.005_{\pm.000}}$.$000_{\pm.001}.$.$001_{\pm.004}.$.$001_{\pm.002}.$
CX race .$\bm{.005_{\pm.000}}$.$\bm{\text{-}.002_{\pm.001}}$.$000_{\pm.003}.$.$\text{-}.001_{\pm.002}- .$
In this section, we study the disparate benefits effect of Deep Ensembles using the experimental setup described in Sec. 4. First, we investigate the disparate benefits effect on the FairFace (FF) test dataset. Second, we apply the same models trained on FF to the UTKFace (UTK) dataset. UTK contains facial images similar to those of FF but from a different source, representing a realistic setting for facial analysis under slight distribution shifts. Third, we investigate the disparate benefits effect on the CheXpert (CX) medical imaging dataset to assess whether the impact on fairness of Deep Ensembles also occurs in other domains than facial analysis. Our analysis examines two primary facets of the disparate benefits effect: (i) the relationship between the number of ensemble members and the changes in performance and fairness violations (Fig. 1); and (ii) for which targets and protected group attributes a statistically significant disparate benefits effect is observed (Tab. 1).
Facial analysis (FF). The top row of Fig. 1 shows results for FF, where models were trained on target age and evaluated under the protected group attribute gender. We find that performance increases while fairness decreases when adding ensemble members. In particular, the largest decrease in fairness occurs when the first member is added to the Deep Ensemble. Tab. 1 presents the change ($\Delta$) in performance and fairness violations between individual models and a Deep Ensemble of 10 members for all tasks. While performance always increases for the Deep Ensemble (positive $\Delta$), fairness does not necessarily increase after ensembling. We observe a disparate benefits effect with significant changes in the fairness metrics for four out of six target / protected group combinations. This occurs primarily when individual members already exhibit substantial fairness violations (gray cells in Tab. 1). The strongest disparate benefits effect (largest absolute $\Delta$) has a negative impact, thus decreasing fairness, i.e. increasing fairness violations. However, there are also cases where the Deep Ensemble is a more fair classifier than individual models (negative $\Delta$).
Facial analysis under a distribution shift (UTK). The middle row of Fig. 1 depicts the results on the UTK dataset, with the same target and protected group as for FF. Individual ensemble members exhibit higher fairness violations than for FF, which can be explained by the distribution shift between FF and UTK. However, the magnitude and behavior of the disparate benefits effect when adding ensemble members are similar to those observed with the FF dataset. The results for all target / group combinations are presented in Tab. 1. The results for UTK are generally similar to those reported for the FF dataset. A notable exception is that the difference in SPD with target variable race and protected group attribute age is of opposite sign and larger for UTK than for FF.
Medical imaging (CX). The bottom row of Fig. 1 shows the results on the CX dataset with age as protected group attribute. The disparate benefits effect also occurs in this task, but with a smaller magnitude, which is explained by the smaller performance gains of Deep Ensembles on this dataset. Similarly as with the facial dataset, the change in fairness after adding the first ensemble member is the most pronounced. The complete results for all protected groups are presented in Tab. 1. For the protected group age, the disparate benefits effect occurs under all fairness metrics. Moreover, there is a significant difference in SPD for the protected group race, although individual models do not have substantial SPD and vice versa for EOD.
Additional results. We investigate the influence of the model size of the individual ensemble members in Apx. F.2. Our results show that for tasks where the disparate benefits effect occurs, it increases with model size. Furthermore, we also analyze the disparate benefits effect under different model architectures. The results and more details are given in Apx. F.3, finding that the results provided in the main paper are consistent across architectures. Finally, we also show the disparate benefits effect for heterogeneous Deep Ensembles in Apx. F.4. Complementary to our main investigation, we explore the notion of minimax fairness (Martinez et al., 2020) within our experiments in Apx. F.1.
Overall, our results demonstrate that Deep Ensembles can decrease fairness. Therefore, we next investigate the reason for their disparate benefits and explore mitigation strategies.
@@ 6 What is the Reason for Disparate Benefits?
In this section, we investigate the potential causes of the disparate benefits effect. The considered fairness metrics (Eq. (5) - Eq. (7)) are based on the per-group PR, TPR and FPR metrics. Therefore, as a first step, we analyze how these change when adding additional ensemble members. Although this analysis offers insight into why the disparate benefits effect occurs, it does not directly explain its underlying cause. We hypothesize that the effect arises from differences in predictive diversity among ensemble members. Our empirical results support this hypothesis, indicating that a gap in the average predictive diversity between groups is a key driver of the effect. To validate this further, we designed two synthetic experiments to investigate this hypothesis in a controlled setting.
@@
Figure 3: Average predictive diversity ($\overline{\text{DIV}}$) per group $A$ and target $Y$. Exemplary results for datasets FF, UTK and CX. Arrows indicate per-target group differences. Top row (a) – (c): Significant disparate benefits (cf. Tab. 1) occur when $\overline{\text{DIV}}$ differences between groups are large. Bottom row (d) – (f): No significant disparate benefits (i.e. equal benefits) occur when $\overline{\text{DIV}}$ differences are small.
@@
Changes to predictions for increasing ensemble size. We begin by examining how the metrics PR, TPR, and FPR for each group change when ensemble members are added, since the considered fairness metrics (Eq. (5) - Eq. (7)) are based on these. Fig. 2 shows these changes for the model trained on FF with age as target variable and gender as protected group, evaluated on the FF test dataset. The results show that the increase in SPD results from a decrease in the PR of the disadvantaged group when adding members to the ensemble, while the PR of the advantaged group remains stable ($\Delta\approx 0$). The TPR of the disadvantaged group stays constant, but the TPR of the advantaged group increases, so the Deep Ensemble improves in correctly predicting $Y=1$ only for the advantaged group, resulting in a higher EOD. The FPR of both groups decreases, more so for the disadvantaged group, thus the Deep Ensemble improves in correctly predicting $Y=0$ (since FPR is one minus the true negative rate). However, this does not offset the TPR disparity, resulting in higher AOD. The results for the remaining tasks are provided in Fig. 12 - 14 in the appendix.
@@
Predictive diversity of ensemble members. The ensemble predictive distribution (Eq. (1)) is an average over the predictive distributions of its members. Therefore, the disparate benefits effect must stem from the characteristics of the predictive distributions of individual members. Previous work investigated the predictive diversity of individual members as the driving mechanism for the increase in the performance of Deep Ensembles (Abe et al., 2022b; Jeffares et al., 2023; Abe et al., 2024). Only if individual members have different predictive distributions, combining them can lead to an ensemble that performs better than the individual models. While previous work investigates predictive diversity for individual inputs $\bm{x}$, we are interested in the average predictive diversity (potentially for a given group) on the test dataset. Following from the definition of predictive diversity by Jeffares et al. (2023) (Theorem 4.3), the average predictive diversity $\overline{\text{DIV}}$ is thus given by
@@
$\overline{\text{DIV}}$ $=\ \frac{1}{K}\sum_{k=1}^{K}\underbrace{\log\left(\frac{1}{N}\sum_{n=1}^{N}p(y=y_{k}\mid\bm{x}_{k},\bm{w}_{n})\right)}_{\text{Ensemble Log-Likelihood}}$ (8)
$-\ \underbrace{\frac{1}{N}\sum_{n=1}^{N}\log p(y=y_{k}\mid\bm{x}_{k},\bm{w}_{n})}_{\text{Average Member Log-Likelihood}}\ ,$
@@
for a test dataset $\mathcal{D}^{\prime}=\{(\bm{x}_{k},y_{k},a_{k})\}_{k=1}^{K}$, and a set of $N$ models with parameters $\{\bm{w}_{n}\}_{n=1}^{N}$. In Apx. C we provide further discussion about the average predictive diversity and how it arises as a natural measure of interest from a Bayesian perspective. Intuitively, the average predictive diversity $\overline{\text{DIV}}$ is a measure of how different individual ensemble members predict. Thus if there is higher $\overline{\text{DIV}}$ for one group, this group has more potential to benefit from ensembling.
@@
Consequently, we hypothesize that differences in the average predictive diversity per group cause the disparate benefits effect. To investigate this hypothesis, we consider two sets of tasks for FF, UTK and CX, respectively: those where the disparate benefits effect occurs and those where it does not occur (cf. Tab. 1). The results are depicted in Fig. 3, showing the average predictive diversity $\overline{\text{DIV}}$ per combination of the target variable $Y$ and the protected group attribute $A$. In agreement with our hypothesis, tasks showing the disparate benefits effect (Fig. 3a-c) have substantial differences in average predictive diversity between groups, while tasks without the effect (Fig. 3d-f) show only minimal differences. Results on all tasks are shown in Fig. 15 - 17 in the appendix.
(a) Inputs per target and group.
(b) Change in performance and fairness for increasing sizes of the Deep Ensemble.
@@
Figure 4: Controlled experiment. (a) Overview of the dataset, showing examplary input for different target / group combinations. (b) The performance (accuracy) increases whereas fairness (1-SPD, 1-EOD, 1-AOD) decreases when adding more members to the ensemble. (c, d) The disparate benefits effect is caused by increased PR and TPR, as well as decreased FPR for the group with higher average predictive diversity $A=1$. For the group with smaller average predictive diversity $A=0$, there are no significant changes in PR, TPR and FPR.
@@
Controlled experiment. To further test our hypothesis of the differences in average predictive diversity between groups causing the disparate benefits effect, we conducted a controlled experiment. We use the FashionMNIST (Xiao et al., 2017) dataset and create a binary classification problem with two targets: “T-shirt/top” ($Y=0$) vs “Shirt” ($Y=1$), and two groups, $A=0$ where the same image of the same target is concatenated twice and $A=1$ where two different images of the same target are concatenated. This is done for both the train and test datasets. An illustration of inputs $\bm{x}$ for both targets and groups is given in Fig. 4(a). Naturally, having an input consisting of two different images ($A=1$) should lead to more diverse ensemble members, as they may learn to use the top image, the bottom image or any combination of features from both. The concatenation of two identical images ($A=0$) does not provide additional information and therefore should not lead to an increase in diversity of the ensemble members. This intuition is experimentally confirmed by having a higher $\overline{\text{DIV}}$ for $A=1$ (Fig. 4c). We observe the same behavior regarding the change in performance, fairness (Fig. 4b) and PR, TPR and FPR (Fig. 4d) as for the real-world datasets that we investigate throughout the rest of the paper. In sum, the synthetic dataset (Fig. 4(a)) enforces more predictive diversity for one group (Fig. 4c), leading to the disparate benefits effect (Fig. 4b, d).
Controlling for predictive diversity. Next, we want to alter the level of predictive diversity and further investigate its relationship with the observed changes in fairness metrics. We used a similar setup as in the previous controlled experiment, i.e., the FashionMNIST dataset and the same targets. However, we used a different methodology to define the groups, in order to vary the level of predictive diversity in the advantaged group $A=1$. We define an input $\bm{x}$ for the disadvantaged group $A=0$ as the original image concatenated with uniform random noise of the same size (each pixel is drawn independently). Furthermore, we define an input for the advantaged group $A=1$ as the original image concatenated with a linear interpolation between a different image of the same target and uniform random noise. The linear interpolation coefficient is $\alpha$, where $\alpha=0$ results in solely uniform random noise ($A=0$ and $A=1$ are equivalent then) and $\alpha=1$ results in two concatenated images from the same label. Thus, for $\alpha=1$, $A=1$ is equivalent to how it was defined in the previous controlled experiment. An illustration of inputs $\bm{x}$ for both targets and groups under different values of $\alpha$ is given in Fig. 6.
The results are shown in Fig. 6. In order to summarize the average predictive diversity into a single number, we calculate a diversity score as $|\overline{\text{DIV}}_{Y=1,A=1}-\overline{\text{DIV}}_{Y=1,A=0}|+|\overline{\text{DIV}}_{Y=0,A=1}-\overline{\text{DIV}}_{Y=0,A=0}|$. Intuitively speaking, this is the sum of the lengths of the arrows in Fig. 3, 4c (also in Fig. 15, 16, 17), shown in the rightmost plot in Fig. 6. We observe that for increasing $\alpha$, the diversity score increases. Furthermore, we find that the changes ($\Delta$) in accuracy, SPD, EOD and AOD due to ensembling increase as well, being highly correlated with the average predictive diversity. We provide the absolute accuracies, SPDs, EODs and AODs for individual ensemble members, the Deep Ensemble and the differences between those in Tab. 2 in the Appendix. In sum, for this second controlled experiment, we find that the higher the predictive diversity per group, the stronger the disparate benefits effect.
(a) $\alpha=0$
(b) $\alpha=0.2$
(c) $\alpha=0.4$
(d) $\alpha=1$
Figure 5: Controlling for predictive diversity. Inputs per target $Y$ and group $A$ for different levels of linear interpolation factor $\alpha$.
Figure 6: Change ($\Delta$) in accuracy, SPD, EOD and AOD due to ensembling, as well as the diversity score for different levels of linear interpolation factor $\alpha$. The disparate benefits effect is stronger for experimental conditions with higher average predictive diversity.
7 Mitigating the Unfairness caused by the Disparate Benefits Effect